OpenStudy (anonymous):

A stopper tied to the end of a string is swung in a horizontal circle. If the mass of the stopper is 13.0g, and the string is 93.0cm, and the stopper revolves at a constant speed 10 times in 11.8s. (a) What is the tension on the string?

OpenStudy (anonymous):

v = 4.95 m/s a= 26.4 m/s2 F = 3.4N

OpenStudy (anonymous):

What are these numbers for?

OpenStudy (anonymous):

that's the answer sweetie.

OpenStudy (anonymous):

F=mv2/r

OpenStudy (anonymous):

Force is the tension in the string.